# MMHMR: Generative Masked Modeling for Hand Mesh Recovery

### v1

@@ Abstract
Reconstructing a 3D hand mesh from a single RGB image is challenging due to complex articulations, self-occlusions, and depth ambiguities. Traditional discriminative methods, which learn a deterministic mapping from a 2D image to a single 3D mesh, often struggle with the inherent ambiguities in 2D-to-3D mapping. To address this challenge, we propose MMHMR, a novel generative masked model for hand mesh recovery that synthesizes plausible 3D hand meshes by learning and sampling from the probabilistic distribution of the ambiguous 2D-to-3D mapping process. MMHMR consists of two key components: (1) a VQ-MANO, which encodes 3D hand articulations as discrete pose tokens in a latent space, and (2) a Context-Guided Masked Transformer that randomly masks out pose tokens and learns their joint distribution, conditioned on corrupted token sequence, image context, and 2D pose cues. This learned distribution facilitates confidence-guided sampling during inference, producing mesh reconstructions with low uncertainty and high precision. Extensive evaluations on benchmark and real-world datasets demonstrate that MMHMR achieves state-of-the-art accuracy, robustness, and realism in 3D hand mesh reconstruction. Project website: https://m-usamasaleem.github.io/publication/MMHMR/mmhmr.html.
@@ 1 Introduction
To overcome these limitations, we introduce MMHMR, a novel generative masked model designed for accurate 3D hand mesh recovery. By learning and sampling from the underlying joint distribution of hand articulations and image features, our model synthesizes most probable 3D hand meshes, mitigating ambiguities inherent in single-view reconstruction. MMHMR consists of two main components: a VQ-MANO and a context-guided masked transformer, which are trained in two consecutive stages. In the first stage, VQ-MANO is trained with Vector Quantized Variational Autoencoders (VQ-VAE) [56] to encode continuous hand poses (e.g., joint rotations) of the MONO parametric hand model into a sequence of discrete pose tokens. In the second stage, the token sequence is partially masked and the context-guided masked transformer is trained to reconstruct the masked tokens by learning token conditional distribution, based on multiple contextual clues, including corrupted token sequence, image features, and 2D pose structures.
This generative masked training allows MMHMR to learn an explicit probabilistic mapping from 2D images to plausible 3D hand meshes. Such probabilistic mapping enables confidence-guided sampling during inference, where only pose tokens with high prediction confidence are retained. This sampling process enables MMHMR to leverage measurable uncertainty to mitigate ambiguities in the 2D-to-3D mapping, resulting in enhanced mesh reconstruction accuracy. Our contributions are summarized as follows
  • MMHMR is the first to leverage generative masked modeling for reconstructing robust 3D hand mesh. The main idea is to explicitly learn the 2D-to-3D probabilistic mapping and synthesize high-confidence, plausible 3D meshes by sampling from learned probabilistic distribution.
@@
  • We demonstrate through extensive experiments that MMHMR outperforms SOTA methods on standard datasets.
@@ 2.2 Generative Methods
Our MMHMR employs a generative method that learns a probabilistic mapping from the input image to the output mesh. It utilizes this learned distribution to synthesize high-confidence, plausible 3D hand meshes based on 2D visual contexts. HHMR [37] is the only other generative hand mesh recovery method in the literature. Unlike HHMR [37] that utilizes diffusion models, MMHMR is inspired by the success of masked image and language models for image and text generation tasks [16, 69, 7, 17, 8]. This fundamental difference allows MMHMR to explicitly and quantitatively estimate confidence levels or prediction probabilities for all mesh reconstruction hypotheses, enabling confidence-guided hypothesis selection for accurate reconstruction. In contrast, HHMR’s denoising diffusion process synthesizes multiple mesh hypotheses without associating a confidence level with each hypothesis. Thus, it only reports the theoretically best mesh reconstruction by finding the hypothesis with minimal reconstruction errors under the assumption that ground-truth meshes are available.
3 Proposed Method: MMHMR
@@
Problem Formulation. Given a single-hand image $I$, we aim to learn a mapping function $f(I)=\{\theta,\beta,\pi\}$ that regresses the MANO [49] model parameters from the input image. This mapping function encompasses three key components: the hand pose parameters $\theta\in\mathbb{R}^{48}$, shape parameters $\beta\in\mathbb{R}^{10}$, and camera parameters $\pi\in\mathbb{R}^{3}$, enabling comprehensive 3D hand reconstruction.
@@
Overview of Proposed Method. As depicted in Figure 1, the MMHMR architecture comprises two main modules: VQ-MANO and the Context-Guided Masked Transformer. The process initiates with VQ-MANO, which converts 3D MANO pose parameters ($\theta$) into discrete pose tokens. These tokens are then processed by the Context-Guided Masked Transformer, which includes a image encoder and masked decoder. The encoder extracts multi-scale image features, which, along with 2D pose guidance and unmasked pose tokens, are fused by masked graph transformer decoder. Within the decoder, Graph-Guided Pose Modeling ensures anatomical coherence by modeling joint dependencies, while the Context-Infused Masked Synthesizer fuses image features and token dependencies to learn the probabilistic reconstruction of pose tokens via differential masked modeling. During inference, the model iteratively refines pose predictions, retaining high-confidence tokens and re-masking those with low confidence, leveraging image semantics, inter-token relationships, and 2D pose guidance to progressively improve accuracy. Finally, the reconstruction is completed as the predicted pose ($\theta$) , shape ($\beta$) and camera parameters ($\pi$) are feed into the MANO hand model.
@@
Figure 1: MMHMR Training Phase. MMHMR consists of two key components: (1) VQ-MANO, which encodes 3D hand poses into a sequence of discrete tokens within a latent space, and (2) a Context-Guided Masked Transformer that models the probabilistic distributions of these tokens, conditioned on the input image, 2D pose cues, and a partially masked token sequence.
Figure 2: MMHMR Inference Phase: Confidence-Guided Iterative Sampling — a step-by-step refinement of pose selection by probabilistically sampling high-confidence tokens.
@@ 3.1 Hand Model and VQ-MANO
Our approach utilizes the MANO hand model [49], which takes pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$ as input. The function $M(\theta,\beta)$ outputs a 3D hand mesh $M\in\mathbb{R}^{V\times 3}$ with $V=778$ vertices and joint locations $X\in\mathbb{R}^{K\times 3}$ with $K=21$ joints, enabling both surface and pose representation.
The VQ-MANO is a MANO hand tokenizer that learn a discrete latent space for 3D pose parameters $\theta\in\mathbb{R}^{48}$ by quantizing the continuous pose embeddings into a learned codebook $C$ with discrete code entries, as depicted in Figure 1(a). To this end, we employ a Vector Quantized Variational Autoencoder (VQ-VAE) [56] for pretraining the tokenizer. Specifically, we input the MANO pose parameters $\theta$ into a convolutional encoder $E$, which maps them to a latent embedding $z$. Each embedding $z_{i}$ is then quantized to its nearest codebook entry $c_{k}\in C$ based on Euclidean distance, defined as
@@ 3.2.1 Multi-scale Image Encoder
Our encoder uses a vision transformer (ViT-H/16) to extract image features [46], processing 16x16 pixel patches into feature tokens. Following ViTDet [2], we adopt a multi-scale feature approach by upsampling the initial feature map to produce feature maps at varying resolutions. This multi-scale representation is critical for handling complex articulations and self-occlusions in hand poses. High-resolution maps provide fine-grained joint details, while low-resolution maps capture global hand structure, balancing precision in joint positioning with overall anatomical coherence. Moreover, we utilize cross-attention with low-resolution feature maps in the x-Attention head to regress stable shape parameters ($\beta$) and camera orientation ($\pi$), making the process computationally efficient. This approach decouples shape estimation from pose modeling, preserving morphological stability and spatial alignment, and enhancing robustness and anatomical accuracy in 3D hand reconstruction.
@@ 3.2.2 Masked Graph Transformer Decoder
The Masked Transformer Decoder is composed of two key components: Graph-Guided Pose Modeling (GGPM) and the Context-Infused Masked Synthesizer Module.
Graph-Guided Pose Modeling (GGPM). Our decoder employs 2 blocks of lightweight graph transformer that processes pose tokens generated by VQ-MANO, enriched with 2D pose guidance, where hand pose tokens are represented as graph nodes linked by learnable adjacency matrices to capture joint relationships effectively. To enhance stability and anatomical accuracy, we integrate a transformer encoder with a Squeeze-and-Excitation (SE) block [29], which emphasizes joint orientations and angles and ensures spatial alignment through a 1x1 convolution layer. Within the transformer encoder, Multi-Head Attention (MHA) and pointwise convolution layers refine high-resolution dependencies between joints, producing a cohesive and anatomically aligned 3D pose representation, denoted as $Q_{\text{GGPM}}$. The output $Q_{\text{GGPM}}$ is formulated as:
@@
$Q_{\text{GGPM}}=\text{SE}(\text{Norm}(Q^{\prime}))$ (2)
where $Q_{\text{C}}\in\mathbb{R}^{K\times D}$represents refined MANO pose tokens and 2D pose guidance queries, and $Q_{\text{GGPM}}$ is the stabilized, anatomically consistent pose representation. The refined pose token queries $Q_{\text{GGPM}}$ are then passed to the Context-Infused Masked Synthesize to synthesize a precise 3D hand mesh that integrates global and local anatomical features.
Context-Infused Masked Synthesizer. We leverage a multi-layer transformer whose inputs are refined pose tokens $Q_{\text{GGPM}}$ and cross-attends them with multi-scale feature maps generated by the image encoder. To enhance computational efficiency with high-resolution feature maps, a deformable cross-attention mechanism is employed [72]. This allows each pose token to focus on a selected set of sampling points around a learnable reference point, rather than the entire feature map. By concentrating attention on relevant areas, the model achieves a balance between computational efficiency and spatial precision, preserving essential information for accurate 3D hand modeling. The deformable cross-attention is defined as:
$\text{MCDA}(Q_{\text{GGPM}},\hat{p}_{y},\{x^{l}\})=\sum_{l,k}A_{lyk}\cdot\mathbf{W}x^{l}(\hat{p}_{y}+\Delta p_{lyk})$
where $Q_{\text{GGPM}}$ are refined manopose token queries, $\hat{p}_{y}$ are learnable reference points, $\Delta p_{lyk}$ are sampling offsets, $\{x^{l}\}$ are multi-scale features, $A_{lyk}$ are attention weights, and $\mathbf{W}$ is a learnable weight matrix. With the inclusion of a [MASK] token in the masked transformer decoder, the module can predict masked pose tokens during training, while also facilitating token generation during inference. This approach allows the [MASK] token to serve as a placeholder for final pose token predictions, supporting robust synthesis of occluded or unobserved hand parts for a coherent 3D hand reconstruction.
@@ 3.3 Training: Differential Masked Modeling
Context-conditioned Masked Modeling. We employ masked modeling to train our model that learns the probabilistic distribution of 3D hand poses, conditioned on multiple contextual cues. Given a sequence of discrete pose tokens $Y=[y_{i}]_{i=1}^{L}$ from the pose tokenizer, where $L$ is the sequence length, we randomly mask out a subset of $m$ tokens with $m=\lceil\gamma(\tau)\cdot L\rceil$, where $\gamma(\tau)$ is a cosine-based masking ratio function. Here, $\tau$ is drawn from a uniform distribution $U(0,1)$, and we adopt the masking function $\gamma(\tau)=\cos\left(\frac{\pi\tau}{2}\right)$, inspired by generative text-to-image modeling strategies [7].
Masked tokens are replaced with learnable [MASK] tokens, forming a corrupted sequence $Y_{\overline{\mathbf{M}}}$ that the model must reconstruct. Each token $y_{i}$ is predicted based on the probabilistic distribution $p(y_{i}|Y_{\overline{\mathbf{M}}},X_{2D},X_{img})$, conditioned on corrupted token sequence $Y_{\overline{\mathbf{M}}}$, 2D pose embedding $X_{2D}$, and image prompt $X_{img}$. This approach enables the model to explicitly account for the uncertainty inherent in mapping 2D observations to a coherent 3D hand mesh. The training objective is to minimize the negative log-likelihood of correctly predicting each pose token in the sequence, formulated as follows:
@@
Expectation-Approximated Differential Sampling. The training objective, $\mathcal{L}_{\text{mask}}$, captures stochastic uncertainty in hand mesh reconstruction, enabling precise estimation of the pose parameter $\theta$ within a structured discrete latent space. Recent studies [46] demonstrate that applying auxiliary 3D and 2D joint losses—measuring alignment between predicted and ground-truth joints in both 3D coordinates and their 2D projections—further refines pose recovery. In generative masked model training, integrating these auxiliary losses requires transforming latent pose tokens into the MANO pose parameter $\theta$, a process that involves non-differentiable probabilistic sampling. To address this, instead of sampling the distribution $p(y_{i}|Y_{\overline{\mathbf{M}}},X_{2D},X_{img})$ to obtain the most probable token from codebook, we implement an expectation-based differential relaxation: instead of directly estimating discrete code indices, the model outputs logits $L$ for each token. These logits undergo a softmax operation, producing the token distribution, which are multiplied by the pretrained codebook, resulting in the mean quantized feature representations $\bar{z}=[\bar{z}_{1},\bar{z}_{2},\dots,\bar{z}_{M}]$:
@@
where $L$ represents the logits matrix, CB denotes the codebook, $M$ is the token count, $K$ specifies the codebook size, and $D$ defines the dimensionality of each codebook entry. This mean token embeddings $\bar{z}$ are feed into the decoder to reconstruct the pose parameter $\theta^{\prime}$. Combined with shape parameters $\beta$ and camera parameters $\pi$, this process reconstructs the 3D hand mesh, enabling the computation of both 3D joint loss and 2D projection loss. Since obtaining $\bar{z}$ is differential, the model can be trained end-to-end using the overall loss function
@@
which combines the masked token prediction loss ($\mathcal{L}_{\text{mask}}$), 3D joint loss ($\mathcal{L}_{\text{3D}}$), 2D projection loss ($\mathcal{L}_{\text{2D}}$), and MANO parameter loss $\mathcal{L}_{\text{MANO}}$. Together, these terms collectively minimize discrepancies in the shape and pose parameters within the MANO space, leading to a precise reconstruction of the 3D hand mesh.

### v2

# MaskHand: Generative Masked Modeling for Robust Hand Mesh Reconstruction in the Wild

@@ Abstract
Reconstructing a 3D hand mesh from a single RGB image is challenging due to complex articulations, self-occlusions, and depth ambiguities. Traditional discriminative methods, which learn a deterministic mapping from a 2D image to a single 3D mesh, often struggle with the inherent ambiguities in 2D-to-3D mapping. To address this challenge, we propose MaskHand, a novel generative masked model for hand mesh recovery that synthesizes plausible 3D hand meshes by learning and sampling from the probabilistic distribution of the ambiguous 2D-to-3D mapping process. MaskHand consists of two key components: (1) a VQ-MANO, which encodes 3D hand articulations as discrete pose tokens in a latent space, and (2) a Context-Guided Masked Transformer that randomly masks out pose tokens and learns their joint distribution, conditioned on corrupted token sequence, image context, and 2D pose cues. This learned distribution facilitates confidence-guided sampling during inference, producing mesh reconstructions with low uncertainty and high precision. Extensive evaluations on benchmark and real-world datasets demonstrate that MaskHand achieves state-of-the-art accuracy, robustness, and realism in 3D hand mesh reconstruction. Project website: https://m-usamasaleem.github.io/publication/MaskHand/MaskHand.html.
@@ 1 Introduction
Figure 1: MaskHand is a quantifiable probabilistic HMR method that can learn 2D-to-3D mapping distributions and explicitly estimate confidence levels or prediction probabilities for all mesh reconstruction hypotheses
To overcome these limitations, we introduce MaskHand, a novel generative masked model designed for accurate 3D hand mesh recovery. By learning and sampling from the underlying joint distribution of hand articulations and image features, our model synthesizes most probable 3D hand meshes, mitigating ambiguities inherent in single-view reconstruction as shown in Figure 1. MaskHand consists of two main components: a VQ-MANO and a context-guided masked transformer, which are trained in two consecutive stages. In the first stage, VQ-MANO is trained with Vector Quantized Variational Autoencoders (VQ-VAE) [60] to encode continuous hand poses (e.g., joint rotations) of the MONO parametric hand model into a sequence of discrete pose tokens. The discretization in VQ-MANO is crucial for generative masked modeling, allowing the model to learn per-token probabilities conditioned on the input image. In the second stage, the token sequence is partially masked and the context-guided masked transformer is trained to reconstruct the masked tokens by learning token conditional distribution, based on multiple contextual clues, including corrupted token sequence, image features, and 2D pose structures. The learned distribution enables uncertainty quantification at each prediction step, facilitating confidence-guided sampling to iteratively refine mesh reconstructions. By selectively retaining high-confidence tokens during inference, MaskHand minimizes uncertainty in the 2D-to-3D mapping, enabling high-precision hand mesh recovery, even in occluded and ambiguous scenarios.
  • MaskHand is the first to leverage generative masked modeling for reconstructing robust 3D hand mesh. The main idea is to explicitly learn the 2D-to-3D probabilistic mapping and synthesize high-confidence, plausible 3D meshes by sampling from learned distribution.
@@
  • We show that MaskHand is a generic framework that unifies hand-mesh estimation and hand-mesh generation.
  • We demonstrate through extensive experiments that MaskHand outperforms SOTA methods on standard datasets.
@@ 2.2 Generative Methods
Our MaskHand employs a generative method that learns a probabilistic mapping from the input image to the output mesh. It utilizes this learned distribution to synthesize high-confidence, plausible 3D hand meshes based on 2D visual contexts. HHMR [38] is the only other generative hand mesh recovery method in the literature. Unlike HHMR [38] that utilizes diffusion models, MaskHand is inspired by the success of masked image and language models for image and text generation tasks [18, 72, 8, 19, 9]. This fundamental difference allows MaskHand to explicitly and quantitatively estimate confidence levels or prediction probabilities for all mesh reconstruction hypotheses, enabling confidence-guided hypothesis selection for accurate reconstruction, as shown in Fig. 1. In contrast, HHMR’s denoising diffusion process synthesizes multiple mesh hypotheses without associating a confidence level with each hypothesis. Thus, it only reports the theoretically best mesh reconstruction by finding the hypothesis with minimal reconstruction errors under the assumption that ground-truth meshes are available.
3 Proposed Method: MaskHand
@@
Problem Formulation. Given a single-hand image $I$, we aim to learn a mapping function $f(I)=\{\theta,\beta,\pi\}$ that output the MANO [52] model parameters from the input image. This mapping function encompasses three key components: the hand pose parameters $\theta\in\mathbb{R}^{48}$, shape parameters $\beta\in\mathbb{R}^{10}$, and camera parameters $\pi\in\mathbb{R}^{3}$, enabling comprehensive 3D hand reconstruction.
@@
Overview of Proposed Method. As shown in Figure 2, MaskHand consists of two primary modules: VQ-MANO and the Context-Guided Masked Transformer. Initially, VQ-MANO discretizes 3D MANO pose parameters ($\theta$) into discrete tokens, enabling generative masked modeling and uncertainty quantification. Next, the Context-Guided Masked Transformer leverages an encoder-decoder structure: the encoder extracts multi-scale image features, which, combined with 2D pose cues and unmasked tokens, are processed by the masked decoder. The decoder employs Graph-based Anatomical Pose Refinement to capture joint dependencies and a Context-Infused Masked Synthesizer to fuse visual features and token interactions, learning the probabilistic reconstruction via differential masked modeling. During inference, pose predictions are iteratively refined by selectively retaining high-confidence tokens and re-masking uncertain ones, progressively improving accuracy. Finally, the reconstruction is completed as the predicted pose ($\theta$) , shape ($\beta$) and camera parameters ($\pi$) are feed into the MANO hand model.
@@
Figure 2: MaskHand Training Phase. MaskHand consists of two key components: (1) VQ-MANO, which encodes 3D hand poses into a sequence of discrete tokens within a latent space, and (2) a Context-Guided Masked Transformer that models the probabilistic distributions of these tokens, conditioned on the input image, 2D pose cues, and a partially masked token sequence.
Figure 3: Architecture of Graph-based Anatomical Pose Refinement (GAPR) and Context-Infused Masked Synthesizer, illustrating fusion of anatomical pose dependencies and contextual cues for precise mesh reconstruction.
@@ 3.1 Hand Model and VQ-MANO
Our approach employs the MANO model [52], which maps pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ to a 3D hand mesh $M\in\mathbb{R}^{778\times 3}$ and joint locations $X\in\mathbb{R}^{21\times 3}$, representing both hand surface and pose.
VQ-MANO is a MANO hand tokenizer that employs a 1D convolutional autoencoder to learn a discrete latent space for 3D pose parameters $\theta\in\mathbb{R}^{48}$ by quantizing the continuous pose embeddings into a learned codebook $C$ with discrete code entries, as depicted in Figure 2(a). A key feature is its upsampling mechanism, expanding 16 MANO poses into 64 discrete pose tokens, which enhances spatial detail and improves mesh recovery accuracy (Supplementary Section C). To this end, we employ a Vector Quantized Variational Autoencoder (VQ-VAE) [60] for pretraining the tokenizer. Specifically, we input the MANO pose parameters $\theta$ into a convolutional encoder $E$, which maps them to a latent embedding $z$. Each embedding $z_{i}$ is then quantized to its nearest codebook entry $c_{k}\in C$ based on Euclidean distance, defined as
@@ 3.2.1 Multi-scale Image Encoder
Our encoder uses a vision transformer (ViT-H/16) [2] to extract image features [49], processing 16x16 pixel patches into feature tokens. Following ViTDet [39], we adopt a multi-scale feature approach by upsampling the initial feature map to produce feature maps at varying resolutions. This multi-scale representation is critical for handling complex articulations and self-occlusions in hand poses. High-resolution maps provide fine-grained joint details, while low-resolution maps capture global hand structure, balancing precision in joint positioning with overall anatomical coherence. Moreover, we utilize cross-attention with low-resolution feature maps in the x-Attention head to regress stable shape parameters ($\beta$) and camera orientation ($\pi$), making the process computationally efficient. This approach decouples shape estimation from pose modeling, preserving morphological stability and spatial alignment, and enhancing robustness and anatomical accuracy in 3D hand reconstruction.
@@ 3.2.2 Masked Graph Transformer Decoder
The Masked Transformer Decoder is composed of two key components (Figure 2 and 3): Graph-based Anatomical Pose Refinement (GAPR) and the Context-Infused Masked Synthesizer Module.
Graph-based Anatomical Pose Refinement (GAPR). Our decoder employs two blocks of a lightweight graph transformer that process pose tokens generated by VQ-MANO, enriched with 2D pose guidance, ensuring anatomically coherent hand reconstructions as shown in Figure 3. To construct a hand-kinematic aware 2D pose token, we first process OpenPose 2D keypoints [7] through a Graph Convolutional Network (GCN) with a fixed hand skeleton adjacency matrix, capturing explicit kinematic structures. Simultaneously, VQ-MANO pose tokens, which inherently encode implicit and non-linear joint dependencies, are extracted. These two representations are then fused and refined through two graph transformer block, where learnable adjacency matrices adaptively model complex joint relationships, ensuring robust spatial and structural consistency. To further enhance stability and anatomical accuracy, we integrate a transformer encoder with a Squeeze-and-Excitation (SE) block [31], which emphasizes joint orientations and angles while ensuring spatial alignment through a 1x1 convolution layer. Within the transformer encoder, Multi-Head Attention (MHA) and pointwise convolution layers refine high-resolution dependencies between joints, producing a cohesive and anatomically aligned 3D pose representation, denoted as $Q_{\text{GAPR}}$. The output $Q_{\text{GAPR}}$ is formulated as:
@@
$Q_{\text{GAPR}}=\text{SE}(\text{Norm}(Q^{\prime}))$ (2)
where $Q_{\text{C}}\in\mathbb{R}^{K\times D}$ represents MANO pose tokens and 2D pose context queries, and $Q_{\text{GAPR}}$ is the stabilized, anatomically consistent pose representation. These refined tokens $Q_{\text{GAPR}}$ are then processed by the Context-Infused Masked Synthesizer to generate precise 3D hand meshes.
Context-Infused Masked Synthesizer. We leverage a multi-layer transformer (as shown in Figure 3) whose inputs are refined pose tokens $Q_{\text{GAPR}}$ and cross-attends them with multi-scale feature maps generated by the image encoder. To enhance computational efficiency with high-resolution feature maps, a deformable cross-attention mechanism is employed [75]. This allows each pose token to focus on a selected set of sampling points around a learnable reference point, rather than the entire feature map. By concentrating attention on relevant areas, the model achieves a balance between computational efficiency and spatial precision, preserving essential information for accurate 3D hand modeling. The deformable cross-attention is defined as:
$\text{MCDA}(Q_{\text{GAPR}},\hat{p}_{y},\{x^{l}\})=\sum_{l,k}A_{lyk}\cdot\mathbf{W}x^{l}(\hat{p}_{y}+\Delta p_{lyk})$
where $Q_{\text{GAPR}}$ are refined manopose token queries, $\hat{p}_{y}$ are learnable reference points, $\Delta p_{lyk}$ are sampling offsets, $\{x^{l}\}$ are multi-scale features, $A_{lyk}$ are attention weights, and $\mathbf{W}$ is a learnable weight matrix. With the inclusion of a [MASK] token in the masked transformer decoder, the module can predict masked pose tokens during training, while also facilitating token generation during inference. This approach uses the [MASK] token as a placeholder for pose predictions, enabling robust synthesis of occluded or unobserved hand parts for coherent 3D hand reconstruction.
@@ 3.3 Training: Differential Masked Modeling
Context-conditioned Masked Modeling. We employ masked modeling to train our model that learns the probabilistic distribution of 3D hand poses, conditioned on multiple contextual cues. Given a sequence of discrete pose tokens $Y=[y_{i}]_{i=1}^{L}$ from the pose tokenizer, where $L$ is the sequence length, we randomly mask out a subset of $m$ tokens with $m=\lceil\gamma(\tau)\cdot L\rceil$, where $\gamma(\tau)$ is a cosine-based masking ratio function. Here, $\tau$ is drawn from a uniform distribution $U(0,1)$, and we adopt the masking function $\gamma(\tau)=\cos\left(\frac{\pi\tau}{2}\right)$, inspired by generative text-to-image modeling strategies [8]. Masked tokens are replaced with learnable [MASK] tokens, forming a corrupted sequence $Y_{\overline{\mathbf{M}}}$ that the model must reconstruct. Each token $y_{i}$ is predicted based on the probabilistic distribution $p(y_{i}|Y_{\overline{\mathbf{M}}},X_{2D},X_{img})$, conditioned on corrupted token sequence $Y_{\overline{\mathbf{M}}}$, 2D pose embedding $X_{2D}$, and image prompt $X_{img}$. This approach enables the model to explicitly account for the uncertainty inherent in mapping 2D observations to a coherent 3D hand mesh. The training objective is to minimize the negative log-likelihood of correctly predicting each pose token in the sequence, formulated as follows:
@@
Figure 4: MaskHand Inference Phase: Confidence-Guided Iterative Sampling — a step-by-step refinement of pose selection by probabilistically sampling high-confidence tokens.
Figure 5: Comparison of SOTA methods: MaskHand synthesizes unobserved parts for accurate 3D reconstructions in occluded scenarios.
Expectation-Approximated Differential Sampling. The training objective $\mathcal{L}_{\text{mask}}$ captures stochastic uncertainty in hand mesh reconstruction, enabling precise estimation of pose parameters $\theta$ within a structured latent space. Recent studies [49] show that auxiliary 3D and 2D joint losses improve pose recovery by aligning predicted and ground-truth 2D/3D joints. Integrating these losses requires transforming latent pose tokens into $\theta$, a process involving non-differentiable probabilistic sampling. To address this, instead of sampling the distribution $p(y_{i}|Y_{\overline{\mathbf{M}}},X_{2D},X_{img})$ to obtain the most probable token from codebook, we implement an expectation-based differential relaxation: instead of directly estimating discrete code indices, the model outputs logits $L$ for each token. These logits undergo a softmax operation, producing the token distribution, which are multiplied by the pretrained codebook, resulting in the mean quantized feature representations $\bar{z}=[\bar{z}_{1},\bar{z}_{2},\dots,\bar{z}_{M}]$:
@@
where $L$ represents the logits matrix, CB denotes the codebook, $M$ is the token count, $K$ specifies the codebook size, and $D$ defines the dimensionality of each codebook entry. This mean token embeddings $\bar{z}$ are feed into the decoder to reconstruct the pose parameter $\theta^{\prime}$. Combined with shape parameters $\beta$ and camera parameters $\pi$, this process reconstructs the 3D hand mesh, enabling the computation of both 3D joint loss and 2D projection loss. Since obtaining $\bar{z}$ is differential, the model can be trained end-to-end using the overall loss function .
@@
which combines the masked token prediction loss ($\mathcal{L}_{\text{mask}}$), 3D joint loss ($\mathcal{L}_{\text{3D}}$), 2D projection loss ($\mathcal{L}_{\text{2D}}$), and MANO parameter loss $\mathcal{L}_{\text{MANO}}$. Together, these terms collectively minimize discrepancies within MANO space, enabling precise 3D hand reconstruction.